let f(x)=2x^2 and evaluate (f(x)-f(a))/(x-a)
Answer by tommyt3rd(5050)

(f(x)-f(a))/(x-a)=
(2x^2-2a^2)/(x-a)=
2((x+a)(x-a))/(x-a)=
2x+2a